OpenStudy (anonymous):

Simplify this expression: 7x^2+35x+28/2x^2+13x+20

OpenStudy (anonymous):

\[\frac{7 x^2+35 x+28}{2 x^2+13 x+20}=\frac{(7 (x+1) (x+4))}{((x+4) (2 x+5))}=\frac{7 (x+1) }{ (2 x+5)} \]